Mathematics - Bayes Formula

Jim Orson’s factory obtains​ capacitors from three different​ partners: 40​% from​ A, 10​% from​ B, and 50​% from C.

The defective rates for these partners are 3%, 3%​, and 4​% respectively.

If a defective​ capacitor is returned, what is the probability that it came from partner A? From​ B? From​ C?
Answer #1

P(Defective) =P(A)*P(defective|A)+P(B)*P(defective|B)+P(C)*P(defective|C)

=0.4*0.03+0.1*0.03+0.5*0.04 =0.035

therefore posterior probability if a defective is found"

P(A|defective )=P(A)*P(defective|A)/P(Defective) =0.4*0.03/0.035=0.3429

P(B|defective )=P(B)*P(defective|A)/P(Defective) =0.1*0.03/0.035=0.0857

P(C|defective )=P(C)*P(defective|A)/P(Defective) =0.5*0.04/0.035=0.5714